OpenStudy (anonymous):

MEDALLLLLLLLL + FAN Which of the values for x and y make the equation 2x + 3y + 4 = 15 true? x = 6, y = 3 x = 3, y = 5 x = 6, y = 4 x = 1, y = 3

OpenStudy (anonymous):

@ArkGoLucky

OpenStudy (anonymous):

Well, I don't think there is any trick you can use here. You just have to plug in the answer choices and see what works

OpenStudy (anonymous):

You could solve for one of the variables in terms of the other like

OpenStudy (anonymous):

\[2x+3y+4=15\]\[2x+3y=11\]\[2x=11-3y\]\[x=11/2 -3y/2\]But that won't really help